the side of a rectangle are 5 and 12. if we inrease both sides by the same amount, the new area is 60 more than the original area, by how much are the sides increased.
Answer by ewatrrr(24785)

Hi,
Let x represent the amount each side is increased.
area ofa rectangle is = l*w
.
question states
(12 + x)(5 + x) = 5*12 + 60 = 120
.
solve
60 + 17 x + x^2 = 120
.
x^2 + 17x - 60 = 0
.
factor
(x + 20)(x-3) = 0
.
cannot use solution x = -20
.
x = 3
.
check your answer
15 * 8 = 120